OpenStudy (anonymous):

The function H(t) = -16t2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. Part B: What is the maximum height that the projectile will reach? Show your work. Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in fe

OpenStudy (anonymous):

the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know?

OpenStudy (anonymous):

@perl @perl

OpenStudy (perl):

Which part would you like help with?

OpenStudy (anonymous):

all of it.....

OpenStudy (perl):

Ok we can start with part a)

OpenStudy (anonymous):

maybe once i get started i MIGHT be able to do it thanks so much

OpenStudy (anonymous):

H(t) = -16t^2 + 80t + 96

OpenStudy (anonymous):

i think ^^

OpenStudy (perl):

We are given that \( H(t) = -16t^2 + vt + s\) shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A) The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second. That means s = 96 and v = 80

OpenStudy (perl):

yes that's correct :)

OpenStudy (anonymous):

ok so my equation becomes : H(t) = -16t^2 + 80t + 96

OpenStudy (anonymous):

ok so thats part A right?

OpenStudy (anonymous):

now part b i have to find the maximum

OpenStudy (anonymous):

so the axis of symmetry is 2.5?

OpenStudy (perl):

the maximum or minimum of a parabola occurs at \(\large t= \frac { -b}{2a} \)

OpenStudy (anonymous):

ok so I got a maximum of 196

OpenStudy (perl):

how plug in that time t = 2.5 to find the maximum height

OpenStudy (anonymous):

196 :)

OpenStudy (perl):

correct

OpenStudy (anonymous):

ok so now part c... now i really dont know where to start

OpenStudy (perl):

Part C) the directions are not complete . Can you check

OpenStudy (anonymous):

yes hang on one sec please

OpenStudy (anonymous):

Part C: Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet, of the object from the ground at time t seconds.

OpenStudy (anonymous):

Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values]

OpenStudy (perl):

We want to find when \( H(t) = g(t) \) which would ordinarily mean solve the equation \( -16t^2 + 80t + 96= 31 + 32.2t \) but instead of solving for t , here they ask us to make tables for the left side and right side separately, and estimate when they are closest

OpenStudy (perl):



OpenStudy (anonymous):

ok so I would solve each equation both H(t) and g(t) by plugging in the numbers 1-5?

OpenStudy (perl):

we can find y values for both H(t) and g(t) and then estimate where they are closest

OpenStudy (anonymous):

ok so we would do that by plugging in 0 right?

OpenStudy (perl):

yes

OpenStudy (anonymous):

ok so g(t) = 31 + 32.2t , g(t) = 31 + 32.2(0) , g(t) = 31 + 0 , g(t) = 31 H(t) = -16t^2 + 80t + 96 , H(t) = -16(0)^2 + 80(0) + 96 , H(t) = 0 + 0 + 96 , H(t) = 96

OpenStudy (anonymous):

so the y-intercept for H(t) = -16t^2 + 80t + 96 is 96 and the y-intercept for g(t) = 31 + 32.2t is 31

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

can you help me with part d @perl

OpenStudy (anonymous):

equation 1: (4,160) equation 2: (4, 159.8) what does it represent in the context of the problem?

OpenStudy (perl):